OpenStudy (anonymous):

Find the product of (x + 3)^2. x^2 + 6x + 9 x^2 − 6x + 9 x^2 − 9 x^2 + 9

OpenStudy (anonymous):

\[ (x+3)^2=(x+3)(x+3) \]Use the FOIL method. \[\begin{align} \text{F.O.I.L. = First Outside Inside Last}\\ \hline \\ &(ax+b)(cx+d)\\ \ \\ \text{First:}&(\color{blue}{ax}+b)(\color{blue}{cx}+d)\to \color{blue}{ax}\cdot\color{blue}{cx}=\color{blue}{acx^2}\\ \text{Outside:}&(\color{green}{ax}+b)(cx+\color{green}{d})\to \color{green}{ax}\cdot\color{green}{d}=\color{green}{adx}\\ \text{Inside:}&(ax+\color{red}{b})(\color{red}{cx}+d)\to \color{red}{b}\cdot\color{red}{cx}=\color{red}{bcx}\\ \text{Last:}&(ax+\color{orange}{b})(cx+\color{orange}{d})\to \color{orange}{b}\cdot\color{orange}{d}=\color{orange}{bd}\\ \hline \\ (ax+b)(cx+d)&=\color{blue}{acx^2}+\color{green}{adx}+\color{red}{bcx}+\color{orange}{bd} \end{align}\]

OpenStudy (anonymous):

So what's the answer?

OpenStudy (anonymous):

You tell me.

OpenStudy (anonymous):

(x+a)^2 = x^2 + 2ax +a^2. Solve using this formula.